Suppose f is a function that satisfies the equation f (x+y) = f(x)+f(y)+x2y+xy2 for all real real numbers x and y.if limit {f(x)/x} as x →0 is equal to 1, then


a) f(x)>0 for x>0 and f(x)<0 for x<0              2)  f1(0)=1    


3)f11(0)=1  4)f111(x) is a constant function


 

Grade:Upto college level

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Aman V Ranga Reddy,

Ans:-f(x+y)=f(x)+f(y)+x²y+y²x

Differntiating With Respect to X keeping y constant we get, f′(x+y)=f′(x)+2xy+y²...............(1)

Now limit{f(x)/x) as x tends to 0 is=1 Hence f′(0)=1.............(2)

Putting x=0 in eq (1) we get

f′(y)=f′(0)+y²=1+y²

f′(y)=1+y² (which is true as we put y=0 it satifies eq 1)

f′′(y)=2y

f′′′(y)=2

Hence options 2 and 4 are correct

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Aman V Ranga Reddy !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur